In the numbers {1,3,5,7,9} construct the following:


• List of possible sample size of 3 that can be taken from this sets of numbers.


• Sampling distribution of the sample means fornthe size of 3 and the standard error.

1
Expert's answer
2022-03-23T02:24:19-0400

Possible samples of size three

(1,3,5), (1,3,7), (1,3,9), (1,5,7), (1,5,9), (1,7,9), (3,5,9), (3,5,7), (3,7,9), (5,7,9)


Sampling distribution

"P(\mu_3=3.00)=P(\mu_3=3.67)=P(\mu_3=7.00)=0.1"

"P(\mu_3=4.33)=P(\mu_3=5.00)=P(\mu_3=5.67)=0.2"






"\sigma=\sqrt{26.33^2-4.996^2}=25.855"

"S.E.=\frac{25.855}{\sqrt{3}}=14.9274"